OpenStudy (anonymous):

A function is said to have a vertical asymptote wherever the limit on the left or right (or both) is either positive or negative infinity. For example, the function f(x)= \frac{-3(x+2)}{x^2+4x+4} has a vertical asymptote at x=-2. For each of the following limits, enter either 'P' for positive infinity, 'N' for negative infinity, or 'D' when the limit simply does not exist. \displaystyle{ \lim_{x\to -2^-} \frac{-3(x+2)}{x^2+4x+4} = } \displaystyle{ \lim_{x\to -2^+} \frac{-3(x+2)}{x^2+4x+4} =} \displaystyle{ \lim_{x\to -2} \frac{-3(x+2)}{x^2+4x+4} =}

OpenStudy (ash2326):

\[ \lim_{x\to -2^-} \frac{-3(x+2)}{x^2+4x+4} \] \[ \lim_{x\to -2^+} \frac{-3(x+2)}{x^2+4x+4} \] \[\lim_{x\to -2} \frac{-3(x+2)}{x^2+4x+4} \]

OpenStudy (shayaan_mustafa):

these are one sided limits

OpenStudy (anonymous):

thanks@ ash2326

OpenStudy (anonymous):

I DONT UNDERSTAND one sided limits

OpenStudy (anonymous):

i know the limit if undefined at -2

OpenStudy (anonymous):

and thats abt all i know..please help

OpenStudy (ash2326):

\[ \lim_{x\to -2^-} \frac{-3(x+2)}{x^2+4x+4} = \] Let \[x=-2-h\ where\ h\to 0\] so \[ \lim_{h\to 0} \frac{-3(-2-h+2)}{(-2-h)^2+4(-2-h)+4} =\lim_{h \to 0}\frac{-3\times -h}{4+h^2+4h-8-4h+4} \] We get \[\lim_{h\to 0}\frac{3h}{h^2}=+\infty\] @rukh now you try others this way

OpenStudy (anonymous):

why did u use h?

OpenStudy (ash2326):

Whenever we have limit like \(\lim_{x\to2^+}\) it means x is approaching 2 from right side, which means x is just larger than 2, so x can be written as x=2+h, where h is a very small quantity, approaching 0

OpenStudy (ash2326):

Similarly \[\lim_{x\to 2^-}\] means x is approaching 2 from left side, which means x is just smaller than 2 or x=2-h where h is a small quantity tending towards 0

OpenStudy (anonymous):

ok i will try this and see if i get the correct answer

OpenStudy (ash2326):

Yeah you know even if the limits both sides are same. If the two sides limits are infinity and at that point also if it's the same. Then also the limit won't exist

OpenStudy (anonymous):

ok im trying to work it now

OpenStudy (ash2326):

Good

OpenStudy (anonymous):

wait how did u get positive infinity for the example u showed me...when i work it out, i get undefined

OpenStudy (anonymous):

is undefined the same as + infinity????

OpenStudy (ash2326):

If limit is infinity positive or negative or of the form like 0/0, 1^(infinity) or , infinity/infinity. Then the limit doesn't exist

OpenStudy (anonymous):

ok. i see now.

OpenStudy (anonymous):

so would u advise me...for future problems with this same form...to replace x with for example -2+h or -2-h?

OpenStudy (anonymous):

or is there a shorter way to do this?

OpenStudy (ash2326):

Yeah this is the way. No shortcuts:)

OpenStudy (anonymous):

ok well thanks for your help. it was much needed...i was never told to approach the problem this way.

OpenStudy (ash2326):

Welcome @rukh